OpenStudy (anonymous):

Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2.

OpenStudy (anonymous):

Use the slope formula. \[m = \frac{y_1 - y_2}{x_1 - x_2} \implies -2 = \frac{9 - 5}{x - 10} \implies -2(x - 10) = 4\]Can yuo finish it up?

OpenStudy (anonymous):

-2x+20=4 ?

OpenStudy (anonymous):

Yes. Now what do you think you should do?

OpenStudy (anonymous):

-2x = -16 Can you solve for x now?

OpenStudy (anonymous):

=8 . thanks calcmathlete